OpenStudy (anonymous):

ok, so, i need help with this question:, all working please the "give way" sign is in the shape of a equalateral triangle with a vertical height of 41.2 cm a- how long was each side of the road sign? b- what area of metal was used to make the sign

OpenStudy (anonymous):

all you need to know sin = oppsite/hypotenuse

OpenStudy (anonymous):

let the side be 2x now the height will form a right angled triangle with half the base and one side so height ^2 + 1/2 base ^2 = side ^2 so 41.2^2 + x^2 = 4x^2 3x^2 = 1697.44 x^2 = 565.81 x = 23.78 2x = 47.57 the side has length 47.57

OpenStudy (anonymous):

and area = (1/2) absin(C)

OpenStudy (anonymous):

the area = 1/2 *xy * sin theta = 1/2 * 0.86 * 47.57*47.57 =980 cm^2